OpenStudy (anonymous):

find the following quantities exactly. for each, indicate where the quantity shows up on the unit circle. A) sin(5pi/4) A) tan(-pi/3)

OpenStudy (anonymous):

sin(5pi/4) = sin ( pi +pi/4 ) = -sin (pi/4) = -1/ sqrt(2) tan(-pi/3) = -tan(pi/3) = -1/sqrt(3)